You have money in an account at 7% interest, compounded annually. To the nearest year, how long will it take for your money to double?
Answer by Alan3354(69443)

1.07^n = 2
n*log(1.07) = log(2)
n = log(2)/log(1.07)
n =~ 10.244 years
--> 11 years to be ~2x